# Enhancing Variable Selection in Large-scale Logistic Regression: Leveraging Manual Labeling with Beneficial Noise

### v1

@@ II-B1 Beneficial Label Noise
$\tilde{y}_{i}\mid\bm{x}_{i}\sim\text{Binomial}(1,{\Pr}_{1}).$ (10)
@@
The manual labeling model (10) is of particular interest because the joint distribution of the manual label $\tilde{y}_{i}$ and the feature vector $\bm{x}_{i}$ is identical to that of the truth label $y_{i}$ and the feature vector $\bm{x}_{i}$. The joint distributions $(\bm{x}_{i},\tilde{y}_{i})\sim\tilde{f}(\bm{x}_{i},\tilde{y}_{i};\bm{\beta}^{*})$ and $(\bm{x}_{i},{y}_{i})\sim{f}(\bm{x}_{i},{y}_{i};\bm{\beta}^{*})$ are equal, as shown below:
$\tilde{f}(\bm{x}_{i},\tilde{y}_{i};\bm{\beta}^{*})$ $=f(\bm{x}_{i};\bm{\beta}^{*})\tilde{f}(\tilde{y}_{i}|\bm{x}_{i};\bm{\beta}^{*})$ (12)
$=f(\bm{x}_{i};\bm{\beta}^{*}){f}(\tilde{y}_{i}|\bm{x}_{i};\bm{\beta}^{*})$
$={f}(\bm{x}_{i},\tilde{y}_{i};\bm{\beta}^{*}),$
where the second equation holds due to the fact $\tilde{f}(\tilde{y}_{i}|\bm{x}_{i};\bm{\beta}^{*})={f}(\tilde{y}_{i}|\bm{x}_{i};\bm{\beta}^{*})$. This fact is true since the manual labels $\tilde{y}_{i}$ are sampled in accordance with the posterior probabilities of class membership in (10). Consequently, the joint distribution of $(\bm{x}_{i},\tilde{y}_{i})$ is identical to that of $(\bm{x}_{i},{y}_{i})$.
Based on the above discussion, given $n$ independently and identically distributed observations from the manual labeling model, $\{(\bm{x}_{i},\tilde{y}_{i})\}_{i=1}^{n}$, the distribution of the maximum likelihood estimate will be the same as when using a dataset with the truth labels $\{(\bm{x}_{i},{y}_{i})\}_{i=1}^{n}$. It then follows that the maximum likelihood loss of a classifier trained using $\{\tilde{y}_{i}\}_{i=1}^{n}$ will be the same as that of a classifier trained using the truth labels $\{{y}_{i}\}_{i=1}^{n}$. This serves as the theoretical guarantee that label noise can be beneficial when the label noise process is determined by the posterior probabilities of class membership in (LABEL:pop). As a result, when multiple experts provide manual labels, the resulting manual labels can provide more statistical information compared to the truth labels. In the following, we will explain the method for simulating the generation of manual labels from multiple experts. It is worth noting that this generation method is particularly simple and has a low computational burden.
II-B2 Multi-Expert Manual Labeling
@@
$\tilde{\bm{Y}}_{ik}\mid\bm{x}_{i}\sim\text{Multinomial}(1,\bm{\Pr}),i=1,2,\dots,n,$ (13)
$k=1,\ldots,m.$
@@
$\mathbb{S}_{i}\sim\text{Multinomial}(m,\bm{\Pr}).$ (14)
When $m=1$, the $S_{i}^{(1)}$ generated by (14) is the same as the $\tilde{y}_{i}$ generated in (10).
To loosen the assumption regarding the accuracy of manual labeling, we employ a random effects model for the counts $\mathbb{S}_{i}$. The Dirichlet-Multinomial model [30] is an extension of the Multinomial distribution. It can be utilized to take into account the fact that the group of experts ($m>1$) does not possess precise knowledge of the posterior probabilities of $\bm{\Pr}$. We can introduce the overdispersion parameter $\alpha_{0}\in\mathbb{R}^{+}$ to formulate the proposed Dirichlet - Multinomial model for the aggregated noisy manual labels as follows:
$\mathbb{S}_{i}\sim\text{Dirichlet - Multinomial}(m,\alpha_{0}\bm{\Pr}),$ (15)
where $\mathbb{E}(S_{i}^{(1)})=m{\Pr}_{1}$, $\mathbb{E}(S_{i}^{(0)})=m{\Pr}_{2}$, $\text{Var}(S_{i}^{(1)})=m{\Pr}_{1}(1-{\Pr}_{1})\frac{m+\alpha_{0}}{1+\alpha_{0}}$, and $\text{Var}(S_{i}^{(0)})=m{\Pr}_{2}(1-{\Pr}_{2})\frac{m+\alpha_{0}}{1+\alpha_{0}}$. When $m=1$, $S_{i}^{(1)}$ can only take values of either 0 or 1. In this case, $\mathbb{E}(S_{i}^{(1)})={\Pr}_{1}$ and $\text{Var}(S_{i}^{(1)})={\Pr}_{1}(1-{\Pr}_{1})$. That is, for any value of $\alpha_{0}$, $S_{i}^{(1)}\sim\text{Binomial}(1,{\Pr}_{1})$, which is consistent with the $\tilde{y}_{i}$ generated in (10).
The Dirichlet-Multinomial model (15) can also be hierarchically represented as follows:
@@
    That is to say, $\bm{p}_{i}$ will converge to a 0-1 distribution. Consequently, when $\bm{p}_{i}$ takes the values $\{0,1\}$ or $\{1,0\}$, it implies that multiple independent experts will uniformly assign all the manual labels $\tilde{y}_{i}$ to be either 1 or 0. In other words, the value of $S_{i}^{(1)}(S_{i}^{(0)})$ will be either $m(0)$ or $0(m)$
  • When $\alpha_{0}$ approaches infinity ($\bm{p}_{i}$ is nearly equal to $\bm{\Pr}$), the excess variance vanishes, and the Dirichlet-Multinomial(Beta-Multinomial) model (15) converges to the Multinomial distribution (14).
The Dirichlet-Multinomial model not only maintains a relationship between the probability of a labeling error and classification difficulty but also relaxes the assumptions about the accuracy of manual annotations compared to (14). This scenario is analogous to that in Section 2.3 of [11], where we explicitly establish a link between the probability of labeling error and the classification difficulty through the posterior class-probabilities.
III Variable Selection and Relative Efficiency with Manual Labeling
In this section, we first develop the corresponding statistical theories, which are similar to those in Section II-A, under the manual labeling stated in Section II-B. Then, leveraging the derived asymptotic distribution, we further compute the asymptotic relative efficiency of ALASSO-PLR under manually-labeled (with noise) data in comparison to that under truth labels.
@@ III Variable Selection and Relative Efficiency with Manual Labeling
$f(S_{i}^{(1)}$ $\mid\bm{x},\bm{\beta})$ (16)
$=\left({p}(\bm{x},\bm{\beta})\right)^{S_{i}^{(1)}}\left(1-{p}(\bm{x},\bm{\beta})\right)^{m-{S_{i}^{(1)}}}.$
Based on (II-A1) and (6), we can derive the loss function of ALASSO-PLR, denoted as $L(\bm{X},\bm{y},\bm{\beta})$, which is defined as
@@
where $w_{j}=1/|\beta^{{}^{\prime}}_{j}|^{\gamma}$. The negative empirical log-likelihood function using the aggregated noisy manual labels $\bm{S}$ (its density function is in (16)) with ALASSO regularization is
$L(\bm{X},\bm{S},\bm{\beta})=$ $-\sum_{i=1}^{n}\left[S_{i}\bm{\beta}^{\top}\bm{x}_{i}-m\log\left(1+\exp(\bm{\beta}^{\top}\bm{x}_{i})\right)\right]$ (17)
$+$ $\lambda\sum_{j=1}^{d}w_{j}|\beta_{j}|,$
@@
III-A Variable Selection and Asymptotic Normality
Next, analogous to Section II-A, we will derive the statistical theoretical properties of (17) in the presence of label noise. Before presenting statistical theory, we need the following two assumptions,
@@ III-A Variable Selection and Asymptotic Normality
    $|\phi^{\prime\prime\prime}({\bm{\beta}^{\top}\bm{x}})|\leq M(\bm{x})<\infty\quad\text{and}$
    $\mathbb{E}[M(\bm{x})|x_{j}x_{k}x_{l}|]<\infty,\text{for all}1\leq j,k,l\leq d.$
These conditions are conventional and have been widely used to establish the oracle property of PLR in [13, 18, 19]. They are easy to meet. For details, please refer to the discussion in Section 3.2 of [18].
Next, we will give the oracle property of ALASSO-PLR estimator under manual labeling (with noise). The proof is placed in Appendix A-A.
@@
    Asymptotic normality:
[…]
Under the Dirichlet-Binomial model (15), the ALASSO-PLR method using the noisy labels $\tilde{\bm{Y}}$ gives a consistent estimator of the true coefficient $\bm{\beta}^{*}$ with the oracle property. When $m=1$ or $\alpha_{0}\to 0$, $\sqrt{n}\left(\tilde{\bm{\beta}}_{\mathcal{A}}-\bm{\beta}^{*}_{\mathcal{A}}\right)\to N\{0,I_{\mathcal{A}}^{-1}\}$, which is the same as the asymptotic distribution of PLR (under truth labels) in Section II-A2. The following will briefly explain the reasons behind this situation.
@@
  • When $\alpha_{0}$ approaches 0, as discussed in Section II-B2, the manual labels provided by all experts are identical. This situation is equivalent to having a single expert manually assign the labels.
@@
III-B Asymptotic Relative Efficiency
@@
$\text{ARE}=\lim_{n\rightarrow\infty}\frac{\mathbb{E}\{\text{err}(\hat{\bm{\beta}};\bm{\beta}^{*})\}-\text{err}(\bm{\beta}^{*};\bm{\beta}^{*})}{\mathbb{E}\{\text{err}(\tilde{\bm{\beta}};\bm{\beta}^{*})\}-\text{err}(\bm{\beta}^{*};\bm{\beta}^{*})}.$ (18)
@@
Based on Theorem 1, we can obtain the following asymptotic relative efficiency. Its proof is in Appendix A-B.
@@
$\text{ARE}=\frac{m(1+\alpha_{0})}{m+\alpha_{0}}.$ (19)
@@
From the above discussion, it can be concluded that the clearer our understanding of the posterior probability and the larger the number of experts, the higher the asymptotic relative efficiency estimated using the manually-generated labels (with noise) in Section II of this paper. This indicates that such noise is beneficial for variable selection in ALASSO-PLR.
III-C Discussions and Applications
Theorems 1 and 2 reveal that when manual labels are generated according to Section II-B, even if noise occurs, it will not affect the variable selection results and improve asymptotic efficiency of ALASSO-PLR. In Section II-B, we allow the manual labels do not possess precise knowledge of the posterior probabilities of $\bm{\Pr}$. This means that when generating manual labels, we do not necessarily need the true coefficient $\bm{\beta}^{*}$. An appropriate estimate of $\bm{\beta}^{*}$ is sufficient, of course, the more accurate the estimate (the larger the $\alpha_{0}$), the better the performance.
@@ III-C Discussions and Applications
    Enhancing variable selection in large-scale PLR. For the collected data $\{\bm{X},\bm{y}\}$, we can obtain a ordinary PLR estimator $\hat{\bm{\beta}}$ to replace $\bm{\beta}^{*}$ in (13) generating manual labels. The use of these generated labels can enhance the coefficient estimation performance of PLR. The generation of these manual labels is easy, just a multinomial distribution, and does not require too much computational burden.
@@
    Filling in the missing labels for large-scale PLR. When there are missing data in the collected data, an estimated $\hat{\bm{\beta}}$ can be obtained from the existing data and then used to substitute this $\hat{\bm{\beta}}$ into the manual label generation model in (13). The manually generated labels can be employed to fill in the missing labels and improve the coefficient estimation performance of PLR.
IV Partition-Insensitive Parallel Algorithm
@@
Recall the optimization expression in (17). Let $\bm{r}=(r_{1},r_{2},\dots,r_{n})^{\top}=\bm{X}\bm{\beta}=(\bm{\beta}^{\top}\bm{x}_{1},\bm{\beta}^{\top}\bm{x}_{2},\dots,\bm{\beta}^{\top}\bm{x}_{n})^{\top}$. Then, the equivalent constrained optimization form of (17) is written as
$\min_{\bm{\beta},\bm{r}}$ $\sum_{i=1}^{n}\left[-S_{i}r_{i}+m\log\left(1+\exp(r_{i})\right)\right]+\lambda\|\bm{w}\odot\bm{\beta}\|_{1},$
s.t. $r_{i}=\bm{\beta}^{\top}\bm{x}_{i},\quad i=1,2,\dots,n.$ (20)
@@
where $\odot$ denotes Hadamard product. It is easy to see that (IV) is an optimization formula with equality constraints and no coupling terms for the optimization variables, which can be solved by using the alternating direction method of multipliers (ADMM) in [31].
IV-A ADMM for ALASSO-PLR with Manual Labels

### v2

@@ 2.2.1 Useful Label Noise
$\tilde{y}_{i}\mid\bm{x}_{i}\sim\text{Binomial}(1,{\Pr}_{1}).$ (2.8)
@@
The manual labeling model (2.8) is of particular interest because the joint distribution of the manual label $\tilde{y}_{i}$ and the feature vector $\bm{x}_{i}$ is identical to that of the truth label $y_{i}$ and the feature vector $\bm{x}_{i}$. The joint distributions $(\bm{x}_{i},\tilde{y}_{i})\sim\tilde{f}(\bm{x}_{i},\tilde{y}_{i};\bm{\beta}^{*})$ and $(\bm{x}_{i},{y}_{i})\sim{f}(\bm{x}_{i},{y}_{i};\bm{\beta}^{*})$ are equal, as shown below:
$\tilde{f}(\bm{x}_{i},\tilde{y}_{i};\bm{\beta}^{*})=$ $f(\bm{x}_{i};\bm{\beta}^{*})\tilde{f}(\tilde{y}_{i}|\bm{x}_{i};\bm{\beta}^{*})=f(\bm{x}_{i};\bm{\beta}^{*}){f}(\tilde{y}_{i}|\bm{x}_{i};\bm{\beta}^{*})={f}(\bm{x}_{i},\tilde{y}_{i};\bm{\beta}^{*}),$ (2.11)
where the second equation holds due to the fact $\tilde{f}(\tilde{y}_{i}|\bm{x}_{i};\bm{\beta}^{*})={f}(\tilde{y}_{i}|\bm{x}_{i};\bm{\beta}^{*})$. This fact is true since the manual labels $\tilde{y}_{i}$ are sampled in accordance with the posterior probabilities of class membership in (2.8). Consequently, the joint distribution of $(\bm{x}_{i},\tilde{y}_{i})$ is identical to that of $(\bm{x}_{i},{y}_{i})$.
Based on the above discussion, given $n$ independently and identically distributed observations from the manual labeling model, $\{(\bm{x}_{i},\tilde{y}_{i})\}_{i=1}^{n}$, the distribution of the maximum likelihood estimate will be the same as when using a dataset with the truth labels $\{(\bm{x}_{i},{y}_{i})\}_{i=1}^{n}$. It then follows that the maximum likelihood loss of a classifier trained using $\{\tilde{y}_{i}\}_{i=1}^{n}$ will be the same as that of a classifier trained using the truth labels $\{{y}_{i}\}_{i=1}^{n}$. This serves as the theoretical guarantee that label noise can be useful when the label noise process is determined by the posterior probabilities of class membership in (LABEL:pop). As a result, when multiple experts provide manual labels, the resulting manual labels can provide more statistical information compared to the truth labels. In the following, we will explain the method for simulating the generation of manual labels from multiple experts. It is worth noting that this generation method is particularly simple and has a low computational burden.
2.2.2 Multi-Expert Manual Labeling
@@
$\tilde{\bm{Y}}_{ik}\mid\bm{x}_{i}\sim\text{Multinomial}(1,\bm{\Pr}),i=1,2,\dots,n,\ k=1,\ldots,m.$ (2.12)
@@
$\mathbb{S}_{i}\sim\text{Multinomial}(m,\bm{\Pr}).$ (2.13)
When $m=1$, the $S_{i}^{(1)}$ generated by (2.13) is the same as the $\tilde{y}_{i}$ generated in (2.8).
Nevertheless, the aforementioned scenario merely depicts an idealized state, and its realization in practical applications is highly improbable. To loosen the assumption regarding the accuracy of manual labeling, we employ a random effects model for the counts $\mathbb{S}_{i}$. The Dirichlet-Multinomial model [25] is an extension of the Multinomial distribution. It can be utilized to take into account the fact that the group of experts ($m>1$) does not possess precise knowledge of the posterior probabilities of $\bm{\Pr}$. We can introduce the overdispersion parameter $\alpha_{0}\in\mathbb{R}^{+}$ to formulate the proposed Dirichlet - Multinomial model for the aggregated noisy manual labels as follows:
$\mathbb{S}_{i}\sim\text{Dirichlet - Multinomial}(m,\alpha_{0}\bm{\Pr}),$ (2.14)
where $\mathbb{E}(S_{i}^{(1)})=m{\Pr}_{1}$, $\mathbb{E}(S_{i}^{(0)})=m{\Pr}_{2}$, $\text{Var}(S_{i}^{(1)})=m{\Pr}_{1}(1-{\Pr}_{1})\frac{m+\alpha_{0}}{1+\alpha_{0}}$, and $\text{Var}(S_{i}^{(0)})=m{\Pr}_{2}(1-{\Pr}_{2})\frac{m+\alpha_{0}}{1+\alpha_{0}}$. When $m=1$, $S_{i}^{(1)}$ can only take values of either 0 or 1. In this case, $\mathbb{E}(S_{i}^{(1)})={\Pr}_{1}$ and $\text{Var}(S_{i}^{(1)})={\Pr}_{1}(1-{\Pr}_{1})$. That is, for any value of $\alpha_{0}$, $S_{i}^{(1)}\sim\text{Binomial}(1,{\Pr}_{1})$, which is consistent with the $\tilde{y}_{i}$ generated in (2.8).
The Dirichlet-Multinomial model (2.14) can also be hierarchically represented as follows:
@@
    That is to say, $\bm{p}_{i}$ will converge to a 0-1 distribution. Consequently, when $\bm{p}_{i}$ takes the values $\{0,1\}$ or $\{1,0\}$, it implies that multiple independent experts will uniformly assign all the manual labels $\tilde{y}_{i}$ to be either 1 or 0. In other words, the value of $S_{i}^{(1)}(S_{i}^{(0)})$ will be either $m(0)$ or $0(m)$ .
  • 2.
    When $\alpha_{0}$ approaches infinity ($\bm{p}_{i}$ is nearly equal to $\bm{\Pr}$), the excess variance vanishes, and the Dirichlet-Multinomial(Beta-Multinomial) model (2.14) converges to the Multinomial distribution (2.13).
The Dirichlet-Multinomial model not only maintains a relationship between the probability of a labeling error and classification difficulty but also relaxes the assumptions about the accuracy of manual annotations compared to (2.13). This scenario is analogous to that in Section 2.3 of [6], where we explicitly establish a link between the probability of labeling error and the classification difficulty through the posterior class-probabilities.
3 Variable Selection and Relative Efficiency with Manual Labeling
In this section, we first develop the corresponding statistical theories, which are similar to those in Section 2.1, under the manual labeling stated in Section 2.2. Then, leveraging the derived asymptotic distribution, we further compute the asymptotic relative efficiency of ALASSO-PLR under manually-labeled (with noise) data in comparison to that under truth labels.
@@ 3 Variable Selection and Relative Efficiency with Manual Labeling
$f(S_{i}^{(1)}$ $\mid\bm{x},\bm{\beta})=\left({p}(\bm{x},\bm{\beta})\right)^{S_{i}^{(1)}}\left(1-{p}(\bm{x},\bm{\beta})\right)^{m-{S_{i}^{(1)}}}.$ (3.1)
Based on (2.3) and (2.4), we can derive the loss function of ALASSO-PLR, denoted as $L(\bm{X},\bm{y},\bm{\beta})$, which is defined as
@@
where $w_{j}=1/|\beta^{{}^{\prime}}_{j}|^{\gamma}$. The negative empirical log-likelihood function using the aggregated noisy manual labels $\bm{S}$ (its density function is in (3.1)) with ALASSO regularization is
$L(\bm{X},\bm{S},\bm{\beta})=-\sum_{i=1}^{n}\left[S_{i}\bm{\beta}^{\top}\bm{x}_{i}-m\log\left(1+\exp(\bm{\beta}^{\top}\bm{x}_{i})\right)\right]+\lambda\sum_{j=1}^{d}w_{j}|\beta_{j}|,$ (3.2)
@@
3.1 Variable Selection and Asymptotic Normality
Next, analogous to Section 2.1, we will derive the statistical theoretical properties of (3.2) in the presence of label noise. Before presenting statistical theory, we need the following two assumptions,
  • 1.
@@ 3.1 Variable Selection and Asymptotic Normality
    $|\phi^{\prime\prime\prime}({\bm{\beta}^{\top}\bm{x}})|\leq M(\bm{x})<\infty\quad\text{and}\quad\mathbb{E}[M(\bm{x})|x_{j}x_{k}x_{l}|]<\infty,\text{for all}1\leq j,k,l\leq d.$ (3.3)
These conventional conditions have been widely used to establish the oracle property of PLR with convex [8] and non-convex (see [13] and [14]) regularization terms. They are easy to meet. For details, please refer to the discussion in Section 3.2 of [13].
Next, we will give the oracle property of ALASSO-PLR estimator under manual labeling (with noise). The proof is placed in supplementary material 1.1.
@@
    Asymptotic normality: n ( ~β_A - β^*_A ) d N{0, m+ α0m(1+ α0)I_A^-1}.
Under the Dirichlet-Binomial model (2.14), the ALASSO-PLR method using the noisy labels $\tilde{\bm{Y}}$ gives a consistent estimator of the true coefficient $\bm{\beta}^{*}$ with the oracle property. When $m=1$ or $\alpha_{0}\to 0$, $\sqrt{n}\left(\tilde{\bm{\beta}}_{\mathcal{A}}-\bm{\beta}^{*}_{\mathcal{A}}\right)\to N\{0,I_{\mathcal{A}}^{-1}\}$, which is the same as the asymptotic distribution of PLR (under truth labels) in Section 2.1.2. The following will briefly explain the reasons behind this situation.
  • 1.
@@
  • 2.
    When $\alpha_{0}$ approaches 0, as discussed in Section 2.2.2, the manual labels provided by all experts are identical. This situation is equivalent to having a single expert manually assign the labels.
@@
3.2 Asymptotic Relative Efficiency
@@
$\text{ARE}=\lim_{n\rightarrow\infty}\frac{\mathbb{E}\{\text{err}(\hat{\bm{\beta}};\bm{\beta}^{*})\}-\text{err}(\bm{\beta}^{*};\bm{\beta}^{*})}{\mathbb{E}\{\text{err}(\tilde{\bm{\beta}};\bm{\beta}^{*})\}-\text{err}(\bm{\beta}^{*};\bm{\beta}^{*})}.$ (3.4)
@@
Based on Theorem 1, we can obtain the following asymptotic relative efficiency. Its proof is in supplementary material 1.2.
@@
$\text{ARE}=\frac{m(1+\alpha_{0})}{m+\alpha_{0}}.$ (3.5)
@@
From the above discussion, it can be concluded that the clearer our understanding of the posterior probability and the larger the number of experts, the higher the asymptotic relative efficiency estimated using the manually-generated labels (with noise) in Section 2 of this paper. This indicates that such noise is useful for variable selection in ALASSO-PLR.
3.3 Discussions and Applications
Theorems 1 and 2 reveal that when manual labels are generated according to Section 2.2, even if noise occurs, it will not affect the variable selection results and improve asymptotic efficiency of ALASSO-PLR. In Section 2.2, we allow the manual labels do not possess precise knowledge of the posterior probabilities of $\bm{\Pr}$. This means that when generating manual labels, we do not necessarily need the true coefficient $\bm{\beta}^{*}$. An appropriate estimate of $\bm{\beta}^{*}$ is sufficient, of course, the more accurate the estimate (the larger the $\alpha_{0}$), the better the performance.
@@ 3.3 Discussions and Applications
    Enhancing variable selection in large-scale PLR. For the collected data $\{\bm{X},\bm{y}\}$, we can obtain a ordinary PLR estimator $\hat{\bm{\beta}}$ to replace $\bm{\beta}^{*}$ in (2.12) generating manual labels. The use of these generated labels can enhance the coefficient estimation performance of PLR. The generation of these manual labels is easy, just a multinomial distribution, and does not require too much computational burden.
@@
    Filling in the missing labels for large-scale PLR. When there are missing or Incomplete data in the collected data [26], an estimated $\hat{\bm{\beta}}$ can be obtained from the existing data and then used to substitute this $\hat{\bm{\beta}}$ into the manual label generation model in (2.12). The manually generated labels can be employed to fill in the missing labels and improve the coefficient estimation performance of PLR.
4 Partition-Insensitive Parallel Algorithm
@@
Recall the optimization expression in (3.2). Let $\bm{r}=(r_{1},r_{2},\dots,r_{n})^{\top}=\bm{X}\bm{\beta}=(\bm{\beta}^{\top}\bm{x}_{1},\bm{\beta}^{\top}\bm{x}_{2},\dots,\bm{\beta}^{\top}\bm{x}_{n})^{\top}$. Then, the equivalent constrained optimization form of (3.2) is written as
$\min_{\bm{\beta},\bm{r}}\$$\sum_{i=1}^{n}\left[-S_{i}r_{i}+m\log\left(1+\exp(r_{i})\right)\right]+\lambda\|\bm{w}\odot\bm{\beta}\|_{1},$
s.t. $r_{i}=\bm{\beta}^{\top}\bm{x}_{i},\quad i=1,2,\dots,n.$ (4.1)
@@
where $\odot$ denotes Hadamard product. It is easy to see that (4) is an optimization formula with equality constraints and no coupling terms for the optimization variables, which can be solved by using the alternating direction method of multipliers (ADMM) in [27].
4.1 ADMM for ALASSO-PLR with Manual Labels